Sol Summary – November 5th

Crew 285 Sol Summary Report 05-11-2023
Sol: 7
Summary Title: Returning to “El Dorado”
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 7 a.m. to have breakfast and prepare the first EVA. At 9 a.m., the first team left for “El Dorado,” a Canyon that was explored and named by a previous Hispanic Crew. We returned to this place to recognize and not forget the legacy of past “Martians”; this team overflew “El Dorado” to take images, create a 3D model for future explorations, and take data for some of our projects. The first EVA team arrived at noon. The crew ate lunch and started preparing the second team that left at 1:30 p.m. back to “El Dorado”.
Look Ahead Plan: Process the data taken during the EVAs
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: Nominal.
EVA: We created a 3D model of the explored Canyon and simulated the rescue protocol with a drone
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report, EVA #5 Report, EVA #6 Report.
Support Requested: Modify the Facebook post, posted on November 3rd, that includes a picture of the Pleiades, which is different from the one our crew astronomer sent. Also, use the name “Hispanics” instead of “Latam” to refer to our crew in media posts.

Operations Report – November 5th

Crew 285 Operations Report 05-11-2023
SOL: 7
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: Yes
Hours: 233.8
Beginning charge: 100%
Ending charge: 51%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 143.5
Beginning charge: 100%
Ending charge: 52%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 256.9
Beginning charge: 100%
Ending charge: 50%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 275.6
Beginning charge: 100%
Ending charge: 47%
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: 23 gallons
Water (static tank): 324 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: no
Summary of internet: No comments
Summary of suits and radios: The battery in suit number 10 died. The fan stopped working after 1 h of use. Once back at the station the battery voltage was measured at 8.5 V with the fan off and 3.0 V with the fan on, while another suit (9) with the same amount of use was measured at 12.6 V off and 12.4 V on. After an hour charging it got to 12 V off but dropped to 8 V on, and the working suit was already at 13.3 V.
Summary of GreenHab operations: relocated the tomato plants and watered
WATER USE: 2.5 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: Not used today
Dual split: Off
Summary of RAM operations: Not used today
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: I request permission to change the battery of suit number 10 given the issue explained above.

Copyright © The Mars Society. All rights reserved. | Main Site